### v3

# CURE: A dataset for Clinical Understanding & Retrieval Evaluation

Abstract
Given the dominance of dense retrievers that do not generalize well beyond their training dataset distributions, domain-specific test sets are essential in evaluating retrieval. There are few test datasets for retrieval systems intended for use by healthcare providers in a point-of-care setting. To fill this gap we have collaborated with medical professionals to create CURE, an ad-hoc retrieval test dataset for passage ranking with 2000 queries spanning 10 medical domains with a monolingual (English) and two cross-lingual (French/Spanish $\rightarrow$ English) conditions. In this paper, we describe how CURE was constructed and provide baseline results to showcase its effectiveness as an evaluation tool. CURE is published with a Creative Commons Attribution Non Commercial 4.0 license and can be accessed on Hugging Face.
CURE: A dataset for Clinical Understanding & Retrieval Evaluation
Nadia Sheikh 1, Anne-Laure Jousse1, Daniel Buades Marcos1, Akintunde Oladipo1, Olivier Rousseau1, Jimmy Lin2 1Clinia, 2University of Waterloo, {nadia.sheikh, anne-laure.jousse, daniel.buades, akintunde.oladipo, olivier.rousseau}@clinia.com jimmylin@uwaterloo.ca
1 Introduction
Evidence-based medicine requires point-of-care healthcare practitioners to keep abreast of advancements in their specialties so they can make data-driven decisions when providing care to patients Sackett (1997). Despite this imperative, it has been estimated that research findings can take up to 17 years to be fully implemented into practice. Balas and Boren (2000).
One impediment for practitioners is the increasingly rapid rate at which relevant literature is published Kamtchum-Tatuene and Zafack (2021). Each year, there are over 1 million new articles published in the field of biomedicine and life sciences González-Márquez et al. (2024).
Given the ever expanding biomedical information landscape, systems that support ad hoc retrieval are an essential tool for health care practitioners to provide evidence-based care Lu (2011). Using ad hoc retrieval, a health-care professional can express their information need as a query, initiating a search that returns likely relevant documents from a vast corpus Lu (2011).
Although the vast majority of published biomedical literature is in English, healthcare providers may practice in other languages Baethge (2008). To effectively accommodate this linguistic diversity, tools aimed at efficient biomedical information access should enable cross-lingual retrieval.
Information retrieval systems and techniques are typically evaluated using test datasets. There are a wide variety of general test datasets available for evaluating retrieval. At its inception, MTEB (Massive Text Embedding Benchmark) made available 15 retrieval test sets, a collection that continues to expand Muennighoff et al. (2022). To the best of our knowledge, there are no cross-lingual datasets suitable for evaluating retrieval systems intended for use by healthcare providers in a point-of-care setting. As such, domain-specific test sets are essential given the dominance of dense retrievers, whose performance does not generalize to out-of-domain tasks Thakur et al. (2021); Ren et al. (2023).
At Clinia, we build products that power search and discovery in health workflows, and the absence of such test datasets has hampered product development, as we cannot benchmark developed solutions systematically. We are working to close this gap by introducing a domain-specific test set which we have already released under an open-source license.
In this work, we introduce Clinical Understanding & Retrieval Evaluation (CURE), a test dataset to evaluate ad hoc retrieval systems for use by medical practitioners in a point-of-care setting. We present results from competitive baselines to demonstrate the utility of CURE as a test dataset and enable comparison with future systems.
CURE is a retrieval dataset with a monolingual and two cross-lingual conditions, with splits spanning ten medical domains. Queries in CURE are natural language questions formulated by health care providers. They express the information needs of practitioners consulting academic literature in the course of their duties. Queries are available in English, French and Spanish. The corpus is constructed by mining an index of english passages extracted from biomedical academic articles. We have made this dataset available in an effort to spur research and development on retrieval in a point-of-care setting.
2 Dataset Construction
CURE is a test collection for passage ranking. In passage ranking, given a corpus C of n passages and a query q, the objective is to retrieve a list of k passages, sorted by decreasing relevance. A relevant passage is one that meets the information need expressed by the query. In the context of a dataset, the relevance of a passage may be marked on a binary (e.g. Relevant or Not Relevant) or graded (e.g. Perfectly Relevant, Partially Relevant, Related, Irrelevant) scale. In CURE passages are marked as Relevant, Partially Relevant and Not Relevant with respect to each query.
CURE includes $20002000$ queries evenly distributed over 10 splits corresponding to the following domains: Dentistry and Oral Health, Dermatology, Gastroenterology, Genetics, Neuroscience & Neurology, Orthopedic Surgery, Otorhinolaryngology, Plastic Surgery, Psychiatry & Psychology and Pulmonology. Each split contains $5050$ layman and $150150$ expert level queries.
In addition to identified relevant passages, CURE’s corpus includes passages mined from an index of over 7777M biomedical passages. These were extracted from articles whose publishers permit their use for evaluation. Sources include the PMC Open Access Subset of Medicine as well as articles scraped from Nature and BioMedCentral. Articles were parsed using the JusText parser and segmented into passages using the parsed document structure. Low-quality passages were filtered out by employing a selection of heuristics introduced to clean training data for Gopher Rae et al. (2021). The corpus contains 244600 passages from 51083 articles. The distribution of passages and articles across domains is presented in Table 5.
CURE was built in three phases: Annotation, Query Translation and Corpus Construction.
During the Annotation phase, medical practitioners expressed their information needs as queries and identified relevant passages in literature the consulted. Annotations were then validated and processed.
During query translation, queries generated by annotators were translated into English(en), French (fr) and Spanish (es) as needed.
The corpus was constructed using relevant passages identified by human annotators and by mining an index of passages extracted from biomedical articles. Large candidate pools were identified for each query and a Large Language Model (LLM) was used to annotate passages for relevance.
After Corpus Construction, the English queries were used to construct the monolingual conditions of CURE while the Spanish and French queries were used to construct the cross-lingual conditions.
2.1 Annotation
To ensure high quality annotations, CURE was constructed with medical professionals. The annotation team included doctors and nurses that were recruited from our professional networks, as well as native speakers of French, English and Spanish. Annotators were compensated for their work. To ensure consistency and quality, annotations were validated and processed. Annotation was carried out over a 9 month period by twelve annotators.
Generation
To create a test dataset with queries representative of the information needs of point-of-care medical practitioners, we asked annotators to formulate queries representing their own information needs and identify relevant passages from biomedical articles they consulted to meet these needs. Annotators were instructed to express queries as natural language questions in English, French or Spanish choosing the language they felt most comfortable with. Annotators identified journal and articles they felt were appropriate given their needs and specialties and we ensured that publishers that permitted their use in evaluation prior to annotation. A copy of our annotation instructions is presented in section 9. An example of a generated question and identified relevant passage is presented in Table 1.
Query: What is the treatment for iatrogenic facial nerve neuroapraxia after facelifting procedure?
Passage: As for the cases of facial nerve neuroapraxia, they were treated by massage of affected muscles, physiotherapies in the form of galvanic stimulation, corticosteroids, and vitamin B12 injection for 8 weeks until complete recovery, which was usually achieved between 3 weeks and 3 months.
Table 1: Example of a generated query and identified relevant passage from the domain of Plastic Surgery
Argilla was used to collect annotations. The interface for annotation is presented in section 9. Annotators provided the text of the query and of identified relevant passages. Additionally, they were asked to specify the query’s language, discipline/s and its level of expertise. Annotators identified queries as being layman or expert with the former being representative of those posed by a medical practitioner seeking to refresh their knowledge or communicate with laymen, and the latter being representative of queries posed by a specialized medical practitioner. An example of a layman and expert query is presented in Table 2.
Layman: What is the most common location of a fracture in facial trauma associated with dental injuries?
Expert: Is there any functionality enabled by Erich arch bars that IMF screws cannot provide?
Table 2: Example of layman query and expert query in the domain of Dental and Oral Health
Validation
Lead annotators from the annotation team validated annotations, ensuring they were of a high quality. In evaluating generated queries, they considered the following criteria:
  • Is the question likely to be asked by a health professional?
    • Does it sound natural?
    • Is it medically relevant?
    • Is it an impactful clinical question?
  • Does the question match the level of expertise it is associated with?
  • Is the question well-written? Considering its syntax, fluency and presence of typos.
  • Is the question specific enough or too broad?
Identified relevant passages were also validated to ensure that they were self-contained and met the information need expressed by the query.
If annotations failed to meet requisite criteria, validators provided feedback and corrected annotations were reevaluated prior to inclusion in the dataset. An example of an invalid annotation with corresponding feedback is presented in Table 3.
Query: What is the impact of physical activity on reducing the risk of gastric cancer, compared to other prevention methods?
Passage: In this condition, seeking prevention methods is key to reducing the burden of gastric cancer. This covers tactics including expanding access to healthcare and screening programs, encouraging healthy lifestyles, and minimizing exposure to established risk factors like smoking and H.pylori infection [35]. One of the easy-accessible methods in this regard is physical activity.
Feedback: The passage is not relevant with respect to the query, or at least, there is no statement on the different impact of physical activity compared to other methods
Table 3: Example of annotation with feedback provided during validation
Processing
Annotators may have entered slight conditions of a single passage as relevant for distinct queries as a passage may meet more more than one information need and the text of relevant passages were entered manually. Furthermore a passage that was identified as relevant for one query may have been embedded in a passage identified as relevant for another query. Although not identified as such, the containing passage would be relevant with respect to the query associated with the contained passage.
In processing annotations, minor variants of the a single passage were identified as such and a representative passage was selected to replace variants. Additionally containing passages were identified and marked as partially relevant for the query associated with the contained passage.
2.2 Query Translation
Using DeepL API, queries were translated so each query is available in English, French and Spanish. DeepL was selected based on its performance relative to alternative tools on a small test of translations evaluated by native speakers. Evaluation criteria included: accuracy, clarity, fluency, appropriateness, consistency, formatting, preservation of proper names and terms, grammar and style, and formal correspondence. After, native speakers performed spot checks to ensure translations were of quality.
2.3 Corpus Construction
Our corpus was constructed by mining for in domain passages and assigning relevance judgments to passages from large candidate pools using an LLM.
Mining
Following Zhang et al. (2023), we mine passages using Reciprocal Rank Fusion (RRF) combining KNN and BM25 retrievers. Iterating through all validated queries, a search was executed using the query and one relevant passage with each type of retriever. Embeddings for the KNN retrievers were generated using BGE-small-v1.5 Xiao et al. (2023). Each KNN retriever considered 250 passages from each shard and returned 50 passages. The BM25 retriever using the query employed a match query with a slop of 2 while the BM25 retriever using the relevant passage employed a match phrase query permitting fuzzy matches. Source articles contributing to the top 50 passages returned were considered to be in-domain and their passages were included in CURE’s corpus.
Candidate Identification
In any sizable retrieval dataset, it is not possible to manually annotate all query-passage pairs for relevance. Typically, small sets of candidates are selected for annotation while the remaining passages are presumed to be irrelevant. As Thakur et al noted, in constructing BEIR, since only small candidate pools are annotated for relevance, datasets can have a bias. For example, Thakur et al. identified a bias in TREC-COVID likely stemming from a use of lexical cues to construct candidate pools Thakur et al. (2021).
To mitigate this bias, we generated large candidate pools for each query using the GTE multi-lingual cross-encoder Zhang et al. (2024). Generated queries and associated passages were assigned to distinct splits, with each split corresponding to one of 10 medical domains. Queries and passages belonging to more than one domain were placed in only one split. Each english query was scored against all passages within the split, and the pool for each query was constructed by taking passages with a score above 0.42. This cut-off score was established using a small test set of passages annotated by humans for relevance and was found to maximize the number of relevant passages included in the pool while minimizing the number of irrelevant passages.
Relevance Judgments
Passages in candidate pools were annotated for relevance using Qwen 2.5 72727272B Yang et al. (2024). The use of LLMs to annotate these pools of passages was inspired by Upadhyay et al.’s work, which found a strong correlation between human relevance judgments and LLM derived judgments Upadhyay et al. (2024). Qwen 2.5 was selected due to its relatively strong concordance with a human on small test set. We used a variation of the prompt used by Upadhyay et al. Generated explanations are included in the dataset. An example explanation is presented in Table 4.
Query: Is treatment of class II malocclusion recommended during childhood?
Passage: The best time and approach to treat Class II malocclusion in one or two stages has been extensively investigated in Orthodontics. Systematic reviews in the literature provide the best scientific evidence, and one of the sources for them is the Cochrane Library . One of these reviews published at Cochrane Library concluded that early treatment of Class II malocclusion and maxillary incisor protrusion in adolescence was not more efficient than treatment in a single stage in the beginning of puberty . Randomized clinical trials conducted in the United States and England , also concluded that, after the second stage of the treatment of Class II malocclusion, there were no differences between the group submitted to early intervention and the group treated in a single stage later. Therefore, the choice of a two- or one-stage treatment seems to be a matter of professional preference, rather than a biological decision.
Judgment: 2
Explanation: The query asks whether treatment of Class II malocclusion is recommended during childhood. The passage provides a detailed review of the literature, including systematic reviews and randomized clinical trials, which conclude that early treatment of Class II malocclusion in childhood (adolescence) is not more efficient than treatment in a single stage during puberty. The passage clearly indicates that the choice between early and late treatment is a matter of professional preference rather than a biological necessity. This information directly addresses the query by suggesting that early treatment is not necessarily recommended over later treatment, providing a clear and actionable answer.
Table 4: Example of a relevance judgment with explanation
After the construction of the corpus, query translations were used to generate a monolingual (English) and two cross-lingual conditions (Spanish-English, French-English) of the dataset with conditions differing only in query language.
3 Dataset Statistics
The average number of tokens per query and passages in the corpus are presented in table Table 5.
The distribution of question types across the dataset is presented in Figure 1 As can be seen, 50.35% of all queries are What questions, followed by 13.45% of queries which are Yes/No questions. In addition to factoid questions, 13.45% of CURE queries are How questions, and 2.9 % are Why questions. On average, each query is associated with 9.93 perfectly relevant passages and 30.43 partially relevant passages.
Figure 1: Query Type Distribution across the domains in CURE
Domain # Passages # Articles # Queries # Judgements Avg. Q Len Avg. P Len
en es fr
Dentistry & Oral Health 29352 5436 200 7966 11.6 13.6 14.6 75.0
Dermatology 15687 3751 200 2692 8.6 10.1 11.2 74.3
Gastroenterology 23068 5832 200 10665 12.0 13.8 14.9 73.9
Genetics 26996 6581 200 13605 11.2 13.9 12.9 87.6
Neuroscience & Neurology 30388 6245 200 8409 12.0 13.9 16.1 79.2
Orthopedic Surgery 24910 4663 200 3002 13.3 15.7 16.6 76.3
Otorhinolaryngology 24192 4424 200 7342 12.4 14.8 15.6 70.7
Plastic Surgery 26605 4995 200 5085 12.6 14.9 16.3 70.6
Psychiatry & Psychology 35756 5967 200 13753 13.5 15.9 16.6 85.1
Pulmonology 32317 5562 200 8197 12.6 15.2 15.8 74.8
244600 51083 2000 80716 12.0 14.1 15.2 77.2
Table 5: Descriptive statics for all domains in CURE. # Avg. Q Len: average number of tokens per query for English (en), Spanish (es), & French (fr); # Avg. P Len: average number of tokens per passage. Tokens are delimited by whitespace.
4 Retrieval Baselines
To demonstrate the effectiveness of our test dataset for retrieval we present results from three different baselines. BM25 serves as our representative for a sparse retrieval baseline. For a dense retrieval baseline we choose a GTE Multilingual Base Zhang et al. (2024) as our representative for models with fewer than 500M parameters and NV Embed V2 as our representative for models with greater than 1B parameters Lee et al. (2024). We will make these baselines available for reproducibility.
Results are presented in Table 6. We used the MTEB evaluation framework Muennighoff et al. (2022) to compute the results. As expected, the the largest dense performance, that of NV-Embed, bests that of the smaller parameter dense model, GTE Multilingual Base, which in turn outstrips the sparse model. Across all retrievers, performance on the monolingual conditions is better than model across cross-lingual conditions of CURE.
NDCG@10 Recall@100
Model Name en-en fr-en es-en en-en fr-en es-en
BM-25 Lù (2024) 0.355 0.018 0.012 0.523 0.044 0.084
mGTE Base (304M) Zhang et al. (2024) 0.558 0.528 0.481 0.714 0.623 0.643
NV Embed v2 (7B) Lee et al. (2024) 0.651 0.603 0.604 0.786 0.738 0.740
Table 6: Sparse and dense baseline results on CURE.
5 Related Test Datasets
In this section, we review widely used biomedical test datasets considering their utility in evaluating ad hoc retrieval systems intended for use by healthcare professionals.
Unlike CURE, many commonly used test datasets that can be leveraged to evaluate retrieval do not reflect the broad range of concerns of medical practitioners. Datasets such as TREC-COVID Roberts et al. (2021) and SciFacts Wadden et al. (2020) have a narrow focus with corpora drawn from CORD-19, a collection of scientific articles on COVID-19 Roberts et al. (2021). The Nutrition Facts Corpus dataset Boteva et al. (2016) is similar with the corpus is constructed from academic texts in the field of nutrition. Other datasets are too broad, extending beyond far the field of medicine. Scifacts Wadden et al. (2020), for example, has a corpus that includes articles drawn from a broad range of scientific disciplines. BioASQ Krithara et al. (2023) as used in BEIR Thakur et al. (2021) has a domain focus most similar to CURE with a corpus constructed from PubMed/MedLine articles. However, queries in the dataset are more representative of the information needs of biomedical researchers than of healthcare providers Krithara et al. (2023).
Dataset Name Query Type Human CLIR Domain #Q Avg Total Corpora
Labels #Labels/Q # Labels Source
Nutrition Facts Boteva et al. (2016) Titles, Descriptions etc. [1] Nutrition 3.8 3237 12334 PubMed Abstracts
TREC-COVID Roberts et al. (2021) Natural questions COVID 50 493.5 24673 CORD-19 Papers
SciDocs Cohan et al. (2020) Article titles [1] Science 1000 29.9 29928
SciFacts Wadden et al. (2020) Assertions COVID 1119 1.1 1258 CORD-19 Papers
BioASQ Krithara et al. (2023) Natural questions Biomedicine 5389 9.7 52413 PubMed
CURE (our work) Natural questions[2] [3] Dentistry & Oral Health, 2000 40.4 80716
Dermatology, Gastroenterology, BioMed Central,
Genetics, Neuroscience & PubMed Central,
Neurology, Orthopedic Surgery, and Nature[1]
Otorhinolaryngology, Plastic Surgery,
Psychiatry & Psychology, and Pulmonology
Table 7: Comparison of related retrieval collections. Human Labels indicates human-generated labels (vs. synthetically generated labels); CLIR indicates whether the collection supports cross-lingual retrieval; Avg # Labels/Q is the average number of labels provided per query; Total # Labels is the total number of human labels (both positive and negative) across all splits. 1Nutrition Facts & SciDocs construct queries from article titles; 2We translate human-generated queries; 3We also annotate query-passage pairs using LLMs as described in subsection 2.3.
Furthermore, many biomedical retrieval datasets are repurposed from datasets originally intended to evaluate systems on downstream tasks other than ranking. As a consequence, queries may not be in forms representative of those posed by actual users. For example, in the Nutrition Facts Corpus dataset Boteva et al. (2016), queries are constructed from article titles and descriptions. The SciDocs dataset Wadden et al. (2020), originally designed to test citation prediction and repurposed for retrieval evaluation in BEIR Wadden et al. (2020), uses paper titles as queries. Similarly, the Scifacts dataset as as used in BEIR Wadden et al. (2020), originally designed to evaluate fact-checking, uses assertions rather than interrogatives as queries. In contrast CURE includes natural language queries formulated by health practitioners.
The relevance judgments in some test retrieval datasets are also limited. For example, in both SciDocs Wadden et al. (2020) and Nutrition Facts Corpus Boteva et al. (2016) relevance is inferred from citations rather than explicitly judged. In CURE, we include explicit human relevance judgments and mitigate bias introduced by relatively shallow candidate pools by using an LLM to fill in missing relevance judgments.
6 Limitations and Future Work
CURE currently covers a limited number of medical domains. As performance in one domain may not be indicative of performance in another domain, in future iterations we would like to expand CURE’s coverage. We would also like to further diversify the type of queries represented in CURE, with a greater focus on queries that require reasoning, particularly within the context of clinical decision making.
Although CURE supports the evaluation of search in academic literature, in a point-of-care setting, a medical practitioner may consult a range of different types of documents, including educational material or Electronic Health Records. In the future, we would like to develop test datasets that enable evaluation of retrieval systems across different types of documents.
7 Conclusion
In this paper we present CURE, a test dataset to evaluate retrieval systems in healthcare settings. CURE was built with a team of healthcare professionals including nurses and doctors, who generated questions and identified relevant passages from biomedical literature. Its corpus was augmented by mining an index of 7777M passages extracted from biomedical articles. To mitigate bias, large candidate pools were constructed and an LLM was used to annotate passages for relevance. The final test dataset consists of 2000 queries evenly distributed over 10 medical domains, and is available in a monolingual (English) and two cross-lingual conditions (French/Spanish $\rightarrow$ English). We also present the performance of three retrievers on this dataset as baselines.
There are few if any test datasets designed to evaluate retrieval systems intended for use by medical practitioners in a point-of-care setting. At Clinia, we use CURE to benchmark our existing retrieval solutions and guide their development. In making CURE freely available for evaluation, we aim to foster research and the development of reliable retrieval tools for healthcare practitioners.
Acknowledgments
We want to highlight our team of annotators, particularly Dr. Arturo Vela Lasagabaster, MSc, MD; Dr. Leila Cattelan, MD; Dr. Natalia Cullell, MD, PhD; and Mr. Noah Tremblay Taillon for their meticulous work. We also thank our CTO, Etienne Soulard-Geoffrion, and the rest of Clinia’s executive team for their belief in responsible ML development and their support for this initiative. Finally, we thank the Governments of Canada and Quebec for their financial support in advancing technological innovation.